OpenStudy (anonymous):

elimination process 5x+5y=-11 7x-3y=19

OpenStudy (saifoo.khan):

x = 1.24 y = -3.44

OpenStudy (angela210793):

5x+5y=-11(*3) 15x+15y=-33 (add both sides) 50x=62 7x-3y=19(*5) 35x-15y=95 x=62/50 y---> 5*62/50 +5y=-11 5y=-11-6.2=-17.2 y=-17.2/5